# Live magnetic observation of parahydrogen hyperpolarization dynamics

### v2

@@ Acknowledgments
The work described is funded by: the Spanish Ministry of Science MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat de Catalunya “Severo Ochoa” Center of Excellence CEX2019-000910-S; the Spanish Ministry of Science projects SAPONARIA (PID2021-123813NB-I00), MARICHAS (PID2021-126059OA-I00), SEE-13-MRI (CPP2022-009771) plus RYC2020-029099-I and RYC2022-035450-I, funded by MCIN/AEI /10.13039/501100011033; Generalitat de Catalunya through the CERCA program; Agència de Gestió d’Ajuts Universitaris i de Recerca Grant Nos. 2017-SGR-1354 and 2021 FI_B_01039; Fundació Privada Cellex; Fundació Mir-Puig; and the BIST–“la Caixa” initiative in Chemical Biology (CHEMBIO). The project has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement 101063517.
The authors are also grateful to Andreas Trabesinger and Sven Bodenstedt for discussions.
Competing interests
The authors declare no competing interests.
Data availability
Experimental data files and processing scripts supporting the findings of this study are provided within the accompanying Supporting Information.

### v3

@@ Acknowledgments
The work described is funded by: the Spanish Ministry of Science MCIN with funding from European Union NextGenerationEU (PRTR-C17.I1) and by Generalitat de Catalunya “Severo Ochoa” Center of Excellence CEX2019-000910-S; the Spanish Ministry of Science projects SAPONARIA (PID2021-123813NB-I00), MARICHAS (PID2021-126059OA-I00), SEE-13-MRI (CPP2022-009771) plus RYC2020-029099-I and RYC2022-035450-I, funded by MCIN/AEI /10.13039/501100011033; Generalitat de Catalunya through the CERCA program; Agència de Gestió d’Ajuts Universitaris i de Recerca Grant Nos. 2017-SGR-1354 and 2021 FI_B_01039; Fundació Privada Cellex; Fundació Mir-Puig; and the BIST–“la Caixa” initiative in Chemical Biology (CHEMBIO); the Helmholtz Association (DB002399). The project has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement 101063517. We are also grateful to Andreas Trabesinger and Sven Bodenstedt for discussions.